MONT 104Q – Mathematical Journeys Third Writing Assignment December, 2015

The third of the larger paper assignments this semester will be devoted to topics connected with our last CHQ common reading of the semester, the memoir Wild by Cheryl Strayed. (Please don’t call it a novel!!)

This one will be slightly different in that the goal is not primarily to defend a particular position on a set question. Instead you will need to delve in detail into a particular passage in the book that represents a key episode or turning point. You should explain why it is a key episode and how it relates to the larger themes of the memoir and/or the CHQ theme. In the process, you should find that you develop a particular point of view about the importance of the passage, and your introductory paragraph should articulate that point of view and prepare the reader to follow what you have to say in the remainder of the essay.

Your writing should show that you have done some careful reading, and some serious thinking about the passage. You may, of course, disagree with ideas presented in class if you want. The point as always is for you to develop your own thinking, and then clearly express it in writing.

Your essay should be 3-5 pages in length (double-spaced, single-sided). Please submit as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Friday, December 11.

For this assignment you have a choice.

1. You may either write on this passage: On page 189, after a phone call with her former husband Paul, Cheryl Strayed says “I missed Paul. I missed my life. But I did not want to go back to it either. That awful moment when Paul and I fell onto the floor and I told him the truth about my infidelities kept coming to me in waves, and I realized that what I’d started when I’d spoken those words hadn’t led only to my divorce but to this: to me sitting alone in Old Station, California, on a picnic table beneath the magnificent sky. I didn’t feel sad or happy. I didn’t feel proud or ashamed. I only felt that in spite of all the things I’d done wrong, in getting myself here, I’d done right.”

2. Or if you prefer, you may select your own passage. The only requirements are that your passage should be roughly one paragraph in length (i.e. about the same length as the passage quoted above) and that it should represent an important moment or turning point in Cheryl Strayed’s understanding of herself, her former life, or the development of her journey. If there is some passage that struck you especially as you read that fits these criteria, by all means “go for it.”

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates if you want. I will also be happy to talk over things and/or read a preliminary draft to give feedback if you like.

2. Especially if you are choosing your own passage, the first paragraph should clearly lay out which passage it is that you will consider and why you think it is important. This will form the thesis and it should be clearly stated and well focused. The reader (me) should not need to guess what your point is.

3. To understand your passage you will probably want to refer to (and/or quote from) other passages from Wild. You should provide interpretations of those passages that come from your own thinking about them. Your evidence should be as specific, detailed, correct, and as varied as possible (focus on small passages and mention the little things that support your point).

4. Do not consult any other sources beyond Wild itself – this means no other books, no web sites, etc.

5. When you use direct quotations, at the end, place page number(s) in parentheses at the end to identify the source.

6. The paper as a whole should be well-organized and “flow” from one point to the next. Your sentences should be clear, readable, and not overly wordy.

7. This paper can be less formal than the two essays from earlier in the semester. You should aim for a lively style and the writing should be in your own voice. (Cheryl Strayed’s voice in her memoir might serve as a model.) But please think twice before using very colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason.

8. Please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

9. Errors in mechanics and usage should not annoy the reader. (As a warning, I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid mistakes in grammar, spelling, punctuation, and word choice.